OpenStudy (anonymous):

What is the equation of the line which includes points (2, 6) and (0, 0)?

OpenStudy (anonymous):

slope is 3 0-6=-6 0-2=-2

OpenStudy (anonymous):

Thank you so much can I ask you for more help? (:

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

umm do yu just need the slope or the line?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

The equation of a line with slope m = -3 and including point (5, 5) is y = -3x + 20.

OpenStudy (anonymous):

True or false

OpenStudy (anonymous):

true if you graph y=-3x+20 5,5 is on the line also y-5=-3(x-5). distribute you get -3x+15 then -5 you get 20

OpenStudy (anonymous):

i meant add 5 to both sides to get 20

OpenStudy (anonymous):

What about this can you help me with this

OpenStudy (anonymous):

true

OpenStudy (anonymous):

This one ?

OpenStudy (anonymous):

you can find any two clear points on the line and then find the slope and that should equal -3. Also the +6 means that on the graph the line should be going through (0,6) and it did so true

OpenStudy (anonymous):

I sent you another one

OpenStudy (anonymous):

umm yes that screenshot right? that is what i just explained

OpenStudy (anonymous):

Here's a new one

OpenStudy (anonymous):

oh i see hold on

OpenStudy (anonymous):

Thank you I really need a good grade on this I really appreciate it (:

OpenStudy (anonymous):

true too

OpenStudy (anonymous):

This one ?

OpenStudy (anonymous):

Would it be the last answer?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

y less than -3x+3

OpenStudy (anonymous):

Can you help me with this one ?

OpenStudy (anonymous):

true

OpenStudy (anonymous):

is this test timed and can you go back?

OpenStudy (anonymous):

Here, and yes I can

OpenStudy (anonymous):

I have an hour let to do 82 more questions

OpenStudy (anonymous):

i would say 2 but im not too confident with that one

OpenStudy (anonymous):

next one

OpenStudy (anonymous):

true

OpenStudy (anonymous):

Next one

OpenStudy (anonymous):

@math&ing001 can you help please?

OpenStudy (anonymous):

this one i dont know wait for that guy and post another on in the mean time

OpenStudy (anonymous):

We could skip it and go to the next question

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

Hold on

OpenStudy (anonymous):

Solve the quadratic equation y2 - 5y = -4 by factoring.

OpenStudy (anonymous):

y=4,1 what are the choices

OpenStudy (anonymous):

that's one of the choices

OpenStudy (anonymous):

ok cool

OpenStudy (anonymous):

next?

OpenStudy (anonymous):

Solve the quadratic equation 5x2 - 125 = 0.

OpenStudy (anonymous):

x=5,-5

OpenStudy (anonymous):

options

OpenStudy (anonymous):

3rd one

OpenStudy (anonymous):

I only have 45 minutes left ! hurry

OpenStudy (anonymous):

its the 3rd on + or - 5!!

OpenStudy (anonymous):

Determine the number of solutions of 2x2 - 3x - 5 = 0.

OpenStudy (anonymous):

2x to the second power

OpenStudy (anonymous):

2

OpenStudy (anonymous):

2 solutions

OpenStudy (anonymous):

The solution of x2 = 15 + 2x is {3, 5}. True or false

OpenStudy (anonymous):

nope -3 and 5

OpenStudy (anonymous):

false

OpenStudy (anonymous):

Help please

OpenStudy (anonymous):

true

OpenStudy (anonymous):

Next one

OpenStudy (anonymous):

b

OpenStudy (anonymous):

Next one

OpenStudy (anonymous):

c

OpenStudy (anonymous):

Here hurrry

OpenStudy (anonymous):

a

OpenStudy (anonymous):

next

OpenStudy (anonymous):

c

OpenStudy (anonymous):

here

OpenStudy (anonymous):

a

OpenStudy (anonymous):

Next one

OpenStudy (anonymous):

c

OpenStudy (anonymous):

Next

OpenStudy (anonymous):

true

OpenStudy (anonymous):

Next

OpenStudy (anonymous):

c

OpenStudy (anonymous):

Next

OpenStudy (anonymous):

hurryyyyy

OpenStudy (anonymous):

a

OpenStudy (anonymous):

plz dont rush me i know ur in a hurry but im literally doing all this for yu hun

OpenStudy (anonymous):

Sorry :(

OpenStudy (anonymous):

its ok :)

OpenStudy (anonymous):

Here next one

OpenStudy (anonymous):

a

OpenStudy (anonymous):

next

OpenStudy (anonymous):

false

OpenStudy (anonymous):

Next

OpenStudy (anonymous):

false

OpenStudy (anonymous):

Here

OpenStudy (anonymous):

a